OpenStudy (anonymous):

show that the following set of functions are subspaces of F(-infinity,infinity) a) all everywhere differentiable function that satisfy f'+2f=0

OpenStudy (saifoo.khan):

Luck>

OpenStudy (anonymous):

thats easy

OpenStudy (anonymous):

all you need are closure under addition scalar multiplication and a zero vector

OpenStudy (anonymous):

and those properties come straight from differenitation

OpenStudy (anonymous):

\[\frac{d}{dx} ( \lambda f(x) ) = \lambda \frac{d}{dx} f(x)\]

OpenStudy (anonymous):

i did not et the point please explain

OpenStudy (anonymous):

\[(f+g)'(x) = f'(x)+g'(x)\]

OpenStudy (anonymous):

and the zero vector (i.e. ) f=0 , obviously solves the equation

OpenStudy (anonymous):

for a subspace you need three things: 1. closure under scalar multiplication 2. closure under vector addition 3. must include a zero vector

OpenStudy (anonymous):

To prove 3, it is obvious , f=0 solves the equation

OpenStudy (anonymous):

but you are not solving according to condition f'+2f=0

OpenStudy (anonymous):

now lets prove 2. consider multiplying both sides by \[\lambda\]

OpenStudy (anonymous):

\[\lambda ( f' + 2f ) = \lambda (0) =0\]

OpenStudy (anonymous):

so that \[(\lambda f )' +2 ( \lambda f ) =0 \]

OpenStudy (anonymous):

which has the exact same form as the original equation , so if f solves the equation then any scalar multiple of f will solve the equation, so its closed under scalar multiplication

OpenStudy (anonymous):

but how under addition

OpenStudy (saifoo.khan):

Sorry, i aint good with this.

OpenStudy (anonymous):

then to prove part one consider two functions f and g that satisify the equations , so that f ' +2f =0 and g' + 2g =0 add these two equations f' +g' + 2f + 2g =0 and remember that the sum of the derivatives is the derivative of the sum, duifferentiation is a linear operation!. group the LHS so: (f+g) ' + 2 (f +g) =0 so if and f and g satisfy it, then their sum will satisfy it , so its closed under vector addition. There it is a subspace.

OpenStudy (anonymous):

remember, you arent solving anything with these questions, its all basic algebra and remember basic things like linearity of differentiation and integration. I did these last year , I can even get some slightly harder questions to put up if you really want , because this is the easiest they will ever be. there is no maths you are just checking those three simple things.

OpenStudy (anonymous):

@elecengineer please send some examples i got ur point thanks

OpenStudy (anonymous):

The solution to the differential equation is of the form \[ce^{-2x}\] Zero vector, c = 0 Additive closure, \[c_2e^{-2x} + c_1e^{-2x} = (c_1+c_2)e^{-2x}\] Multiplicative closure, \[(kc)e^{-2x}\]

OpenStudy (anonymous):

Damn , I thought I thought had some questions involving integrals but turns out I dont

OpenStudy (anonymous):

\[(x+1)p ' (x) - 3p(x) =0 \] show that is a subspace if you want, its pretty much the same ideas , using linear rules of differentiation

OpenStudy (anonymous):

@ sir are you electrical engineer

OpenStudy (anonymous):

or an even more trivial question would be \[p''(x) =0 \] showing thats a subspace is done in 1 line